A 61 block is moving at a velocity of 13.86 m/s on a frictionless surface. A constant force F is applied for 5 seconds which caused the block to accelerate at a rate of 2.86 m/s^2. Find F.


1
Expert's answer
2021-03-22T08:36:27-0400

Assuming mass (m)=61 kg

Acceleration of the block(a)=2.86 m/S2

Therefore force acting on the block is given by,

F=m×a=61×2.86=174.46 N